suppose the mean weight of all the students in this examination hall is 54 kgs, and the standard deviation is 4 kgs. what percentage of students will fall between 44 to 62 kgs? justify your answer


1
Expert's answer
2021-01-14T17:25:29-0500

Let "X" be the weight of student. If the distribution of "X" is not normal we cannot say about the percentage of students will fall between 44 to 62 kgs.

If we assume that the weight of student is normally distributed than "X\sim N (\mu, \sigma^2)".


"Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=54, \sigma=4."


"P(44<X<62)=P(X<62)-P(X\leq 44)"

"=P(Z<\dfrac{62-54}{4})-P(Z\leq\dfrac{44-54}{4})"

"=P(Z<2)-P(Z\leq-2.5)"

"\approx0.97725-0.00621=0.97104"

97.104 % of students will fall between 44 to 62 kgs.